OpenStudy (anonymous):

How many solutions can be found for the equation −4x − 11 = 2(x − 3x) + 13? A. None B. One C. Two D. Infinitely many

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

you can always plug in #'s to see the answer, have you tried that ?

OpenStudy (anonymous):

no @silverr

OpenStudy (anonymous):

try 1 first. −4(1) − 11 = -15. 2(1 − 3(1) + 13, you try this one!

OpenStudy (anonymous):

@poopsiedoodle @AlexandervonHumboldt2

OpenStudy (yanasidlinskiy):

Also you can just use Distributive Property. −4x − 11 = 2(x − 3x) + 13 -4x-11 = 2x - 6x + 13 Your next step is to combine like terms and Simplify. Do you know how to finish off the process?

OpenStudy (anonymous):

no @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Well, as I've said above our next step is to combine like terms on both sides. Do you know what like terms are?

OpenStudy (anonymous):

nah @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Like terms are things such as: 6x + 5x that would be one example. Another one would be 5 +3 and etc. As far as your equation, what are the like terms on one of the sides. Can you tell me?

OpenStudy (anonymous):

@Loser66 @jordanloveangel